OpenStudy (cookiemonster2412):

If you ran 15 km/h for 20 min, how much distance would you cover?

jimthompson5910 (jim_thompson5910):

20 minutes = _____ hours fill in the blank

OpenStudy (cookiemonster2412):

1/3 or 0.333

jimthompson5910 (jim_thompson5910):

Now use the equation D = r*t D = distance r = rate or speed t = time

jimthompson5910 (jim_thompson5910):

r = 15 t = 1/3

OpenStudy (cookiemonster2412):

d=5 km right

jimthompson5910 (jim_thompson5910):

correct